OpenStudy (anonymous):

Kendra has a painting canvas that is 25 inches wide by 31 inches high. She painted a red rose, which covered 40% of the canvas. What was the area of the canvas which was covered by the red rose?

OpenStudy (welshfella):

area of the canvas = 31 * 25 in^2 area covered by red rose = 40% of this = 31*25* 0.40 (note 40% = 0.40 as a decimal)

OpenStudy (anonymous):

would it be 310?

OpenStudy (anonymous):

Tell Kendra that the art industry is dead and she should give up on her dreams and become an accountant

OpenStudy (anonymous):

like her father

OpenStudy (anonymous):

FIRST of all what the hell is on the other 60% of the canvas? What a waste, you could be using it for other reasons but NO. You waste it like all of the other republican LITTERBUGS and let it rot in the ditch.

OpenStudy (anonymous):

guys, i know you're trying to be funny but im looking for a legit answer to the problem.

OpenStudy (anonymous):

find the area and multiply it by .4

OpenStudy (anonymous):

Well trust me you have multiple other problems like illiteracy, another being you are some stupid cat king like psshhh?

OpenStudy (anonymous):

its 320

OpenStudy (anonymous):

no .its .04

OpenStudy (welshfella):

you were correct first time @celestialcatking

OpenStudy (anonymous):

okay, thank you very much for your help!

OpenStudy (welshfella):

yw